OpenStudy (anonymous):

simplify the imaginary number, square root of -17.

OpenStudy (dangerousjesse):

\[\sqrt{-17}\] = 4

OpenStudy (dangerousjesse):

That's just an approximation, though.

zepdrix (zepdrix):

\[\Large\rm \sqrt{-17}=\sqrt{-1\cdot17}=\sqrt{-1}\cdot\sqrt{17}\]We denote the imaginary unit square root of -1 as i, \[\Large\rm \sqrt{-1}=\mathcal i\]Which gives us,\[\Large\rm \mathcal i \sqrt{17}\]

zepdrix (zepdrix):

But yah if you were looking for an approximation.. I guess it's uhhh.. 4-ish 0_o